OpenStudy (anonymous):

evaluate the improper integral, state if it diverges: integral (0 -> infinity) (2dx/25+x^2)

OpenStudy (anonymous):

Ah I know the trick for this one, but first for your understanding: Tell me what the derivative of x^2+25 is, cool? :-)

OpenStudy (anonymous):

2x-25x

OpenStudy (anonymous):

Erm nope :-) How about what the derivative of arctangent is? That's also known as tan\(^{-1}\)( )

OpenStudy (anonymous):

It's just 2x dx for the first one ;-)

OpenStudy (anonymous):

youre right, sorry i confused that with integration, ok so its 2x dx

OpenStudy (anonymous):

arctan is 1/(1+x^2)

OpenStudy (anonymous):

I think what I'm trying to get you to see is two things here... (goes to write) And yes that's correct! gj

OpenStudy (anonymous):

\[\int\limits \frac{2 dx}{25+x^2} = 2 \int\limits \frac{1}{x^2+5^2} dx\]

OpenStudy (anonymous):

scoot that junk out of the numerator so you can see it better

OpenStudy (anonymous):

yeah i see that looks much more beautiful. so its arctan.

OpenStudy (anonymous):

And now for the magic bullet! \[\frac{d}{dx} \frac{1}{x^2+a^2}=\frac{1}{a}\tan^{-1}( \frac{x}{a}) dx\]

OpenStudy (anonymous):

Careful to not forget the 2 outside the integrand though for your final answer. Got it now? :D

OpenStudy (anonymous):

yeah i do so i now have to look at the end behavior of ArcTan as it goes to Infinity

OpenStudy (anonymous):

And that's easy to do if you remember this: inverse switches domain & range So what's the domain for tangent?

OpenStudy (anonymous):

-pi/2 to pi/2

OpenStudy (anonymous):

Yep, guess which is \(\infty\) and which is -\(\infty\)? ;-)

OpenStudy (anonymous):

For arctan

OpenStudy (anonymous):

well the range of tan would be the domain of arc tan correct?

OpenStudy (anonymous):

Yeah the range of tangent is all reall #'s :-)

OpenStudy (anonymous):

The domain of tangent however is restricted

OpenStudy (anonymous):

Remember inverse function are the same as the original in shape except they have been flipped over the line y=x and that means their domains and ranges are swapped compared to the original \(\infty \leftarrow\rightarrow\frac{\pi}{2}\)